If 3.50 ✕ 10−3 kg of gold is deposited on the negative electrode of an electrolytic cell in a period of 2.83 h, what is the current in the cell during that period? Assume the gold ions carry one elementary unit of positive charge. (answer in mA)
Answer #1

3.50 × 10-3 Kg gold = 0.018 mole gold

Total charge in this amount of gold ion is

Q = 0.018× 6.022×1023×1.6×10-19 C

= 1.73 × 102 C

I = Q ÷ t

= { 1.73 ×102 ÷ (2.83×3600)} amp

= 0.01698 amp

= 16.98 mA